OpenStudy (mach_15):

Algebra 2 help please. Really confused

OpenStudy (mach_15):

Consider this function in recursive form f(1) = -2 f(n)= 4f(n-1) where n > select the equivalent function for n >1

OpenStudy (mach_15):

A) f(n) = -2n B) f(n) = -2(n-1) C) f(n) = 2(4)^n-1 D) f(n) = -2(n)^4

satellite73 (satellite73):

\[f(1)=-2\\ f(2)=4f(1)=4\times (-2)\]

satellite73 (satellite73):

\[f(3)=4\times f(2)=4\times 4\times (-2)=-2\times 4^2\]

satellite73 (satellite73):

\[f(4)=4\times f(3)=-2\times 4^3\]

satellite73 (satellite73):

and so on

OpenStudy (mach_15):

I still don't quite get it. I literally have no idea what the question is asking or how to do it.

jimthompson5910 (jim_thompson5910):

`f(1) = -2` means the first term of the sequence is -2 Do you agree with this statement @mach_15 ?

OpenStudy (mach_15):

yeah

jimthompson5910 (jim_thompson5910):

`f(n)= 4*f(n-1)` is the recursive rule it helps you determine the second term, third term, etc to find the second term, plug in n = 2 f(n)= 4*f(n-1) f(2)= 4*f(2-1) ... replace every 'n' with '2' f(2)= 4*f(1) f(2) = 4*(-2) ... replace f(1) with -2 f(2) = -8 So the second term is -8. Notice how i used the first term (f(1) = -2) to find the second term

jimthompson5910 (jim_thompson5910):

let me know if that makes sense or not

OpenStudy (mach_15):

Yes it does

jimthompson5910 (jim_thompson5910):

what is f(3) equal to?

OpenStudy (mach_15):

I don't think I did this right at all but I got 8.

jimthompson5910 (jim_thompson5910):

we'll use the fact that f(2) = -8 let's plug in n = 3 f(n) = 4*f(n-1) f(3) = 4*f(3-1) ... replace every 'n' with '3' f(3) = 4*f(2) f(3) = 4*(-8) ... replace f(2) with -8 (see above) f(3) = -32 make sense?

jimthompson5910 (jim_thompson5910):

if you want to find the fourth term, use the fact that the third term is f(3) = -32

OpenStudy (mach_15):

oh okay I see now. so the fourth term is -128?

jimthompson5910 (jim_thompson5910):

f(n) = 4*f(n-1) f(4) = 4*f(4-1) ... replace every 'n' with '4' f(4) = 4*f(3) f(4) = 4*(-32) ... replace f(3) with -32 (see above) f(4) = -128 yep you got it

jimthompson5910 (jim_thompson5910):

so the sequence of terms is -2, -8, -32, -128, ... and it goes on forever

jimthompson5910 (jim_thompson5910):

The question is: which function in your list of choices also produces that same exact sequence?

jimthompson5910 (jim_thompson5910):

Let's test choice `A) f(n) = -2n` first term happens when n = 1 f(n) = -2n f(1) = -2*1 ... plug in n = 1 f(1) = -2 first term is -2. So far so good second term happens when n = 2 f(n) = -2n f(2) = -2*2 ... plug in n = 2 f(2) = -4 uh oh. We got -4 to be the second term when it should be -8. So choice A can't be the answer because the sequences don't match up Recall that we want -2, -8, -32, -128, ... to be our sequence where n starts at n = 1.

OpenStudy (mach_15):

Oh okay. So I guess ill try B and see if I do it right

jimthompson5910 (jim_thompson5910):

ok tell me what you get

OpenStudy (mach_15):

I got 0. Did I do it right?

OpenStudy (mach_15):

I went ahead and did C and D and C = 2 but D = -2 so the answer would be D right?

jimthompson5910 (jim_thompson5910):

B) f(n) = -2(n-1) first term n = 1 f(n) = -2*(n-1) f(1) = -2*(1-1) f(1) = 0 yep so because the first term (0) doesn't match with the first term we want (-2) this means choice B is out

jimthompson5910 (jim_thompson5910):

D) f(n) = -2(n)^4 let's find the second term using the function in choice D plug in n = 2 f(n) = -2(n)^4 f(2) = -2(2)^4 f(2) = -2*(16) f(2) = -32 we don't have a match. We wanted the second term to be -8 and not -32 So choice D is out

jimthompson5910 (jim_thompson5910):

I'm thinking there's a typo. I think your teacher meant to write -2(4)^(n-1) for choice C

OpenStudy (mach_15):

Maybe. Its an online multiple choice quiz through K12 so I don't know.

OpenStudy (mach_15):

Thank you for all your help! This stuff makes a lot more sense now!

jimthompson5910 (jim_thompson5910):

I'm glad it does